# Compression Hacking: A Supplementary Perspective on Informatics Properties of Language Models from Geometric Distortion

### v2

@@ 4.4 Effect of Anisotropy Razors
The qqplot in Figure 5 illustrates the eigenvalue distributions before and after applying these four razors. The first three methods decouple anisotropy while maintaining the linear geometric structure of the data, resulting in eigenvalues that still exhibit distinct partitioning. In contrast, PCS upscales the low-eigenvalue region, ensuring that the corrected compression relies entirely on the contributions of the principal components. The formal method for anisotropy detection involves examining the “self-normalization” property (i.e., $\mathcal{Z}(\mathbf{c})$ tending toward a constant, independent of $\mathbf{c}$Mu and Viswanath (2018). Figure 5 illustrates the distribution of $\mathcal{Z}(\mathbf{c})$ before and after applying different anisotropy razors. We observe that remove directions leads to a more dispersed $\mathcal{Z}(\mathbf{c})$ distribution, increasing anisotropy. This occurs because truncating certain directions causes the remaining ones to spread more extremely. In contrast, whitening, LW shrinkage, and PCS concentrate the $\mathcal{Z}(\mathbf{c})$ distribution. Notably, PCS achieves more pronounced anisotropy elimination than the other methods by rigidly correcting the eigenvalue distribution.
@@ A.1 Evaluation of Language Models
The evaluation of language models is currently in a state of rapid iterative development, encompassing a variety of tasks, datasets, and benchmarks Celikyilmaz et al. (2020); Zheng et al. (2023); Tan et al. (2024). Traditional evaluation metrics such as accuracy, F1-score Sasaki et al. (2007), BLEU Sellam et al. (2020), and ROUGE Lin (2004) focus on comparing model predictions with annotated labels in downstream tasks. Other metrics like perplexity and cross-entropy loss do not rely on annotated labels and are computed solely based on input text. However, these methods primarily emphasize external evaluation based on model predictions.
Recently, significant efforts have been devoted to exploring the mechanisms by which language models (LMs) process information internally, driving the development of LM self-evaluation Wei et al. (2024); Wang et al. (2024a, b); Zang and Liu (2024a); Dou et al. (2024) independent of specific tasks and model outputs. The concept of “compression as intelligence” has provided an information-theoretic internal evaluation perspective for language models, highlighting that the acquisition of world knowledge by language models is a denoising process Sutskever (2023); Deletang et al. (2023); Wei et al. (2024); Chen et al. (2025). Differential entropy of representations, as a classical information-theoretic measure, effectively quantifies the internal uncertainty of language models Chen et al. (2023a); Zhouyin and Liu (2023). Semantic volume Li et al. (2025) leverages representation-level differential entropy-aware compression metrics to offer a novel perspective for language model evaluation. However, related work has found that such compression can only model the scale of language models and fails to align with their capabilities Zang and Liu (2023b); Wei et al. (2024); Li et al. (2025).
We introduces the concept of compression hacking in language model representations, where the noisy directions of LM representations sacrifice spatial uniformity to feign high compression rates. This implies that we can refine the information compression perspective by considering the geometric distortions in the language model’s representation space.
@@ A.2 Anisotropy of Language Models
Anisotropy The anisotropy of language models reflects the geometric properties of the contextual embedding space. Related studies have observed that during sampling, the spatial embeddings of negative samples exhibit anisotropy, which describes how vectors are distributed within the contextual space Mimno and Thompson (2017); Ethayarajh (2019). The researchers found that most vectors occupy a relatively narrow cone within the space, and that vectors within this cone tend to have high cosine similarity Gao et al. (2019b). Demeter pointed out that using softmax introduces structural weaknesses in the representation space, leading to bias, a common issue in language models Demeter et al. (2020). To better quantify the anisotropy of LMs, related work has identified isolated clusters and low-dimensional manifolds in the contextual embedding space, introducing tools for their qualitative and quantitative analysis Ethayarajh (2019); Cai et al. (2019); Rudman et al. (2022). However, these tools are mainly based on similarity calculations of embedded representations. What is needed instead is an anisotropy metric that can establish a connection with entropy based compression metric.
Anisotropy Razors To mitigate anisotropy in language models, existing research has proposed various solutions. Contrastive learning has emerged as a powerful tool for obtaining effective sentence representations, effectively reducing anisotropy by increasing the spatial distance between positive and negative samples Gao et al. (2021); Zhang et al. (2022b); Jiang et al. (2022); Zang and Liu (2023a, 2024b). In this work, we employ post-processing methods applied directly to the representation space as baseline approaches for the anisotropy razor:
  • Remove Directions Mu and Viswanath (2018): First, subtract the common mean vector of all word vectors to eliminate global bias; then remove the top high-variance principal component directions via Principal Component Analysis (PCA). This process enhances semantic feature discriminability by eliminating non-semantic common information from word vectors, making the word space distribution more isotropic.
  • Whitening Su et al. (2021): Zero-center the representations and transform the covariance matrix into an identity matrix, forcing the embedding distribution toward isotropy.
  • LW Shrinkage Ledoit and Wolf (2004): Linearly shrink the sample covariance matrix toward the diagonal matrices to reduce noise interference in high-dimensional data, yielding more stable covariance matrix estimates. This operation mitigates excessive sensitivity in specific directions, promoting isotropic feature distributions.
@@ Appendix B Statistical Properties of Principal Component Smoothing
Lemma B.1 (Asymptotic Optimality of Ledoit-Wolf Shrinkage Ledoit and Wolf (2004)).
@@
[…]
@@ Appendix C Significance Analysis
Figure 7: The Mann-Whitney U tests of compression-anisotropy regression fitting between different models, where, ***, and **** denote statistical significance at the 0.1%, and 0.01% levels respectively.
@@ Appendix D Implementation Details of the Evaluation Pipeline
Figure 9: The correlation coefficients between compression (PCS) and ground truth under different smoothing coefficient.
For the projection dataset, we primarily collected 1,000 data samples from the pretraining corpus (Wiki Foundation (2025)) and the instruction-tuning dataset (Dolly-15k Conover et al. (2023)) to derive projection data. By sampling thein-context representations of these data points, we aim to estimate the full model’s representation space, ensuring the convergence of our metrics. Our pipeline defaults to sampling 800 data samples. Figure 8 illustrates the cumulative expected values of different metrics as the number of samples increases. We observe that all metrics converge relatively early to stable values, demonstrating that our refined metrics enable robust evaluation based on the provided projection dataset.

### v3

@@ 4.4 Effect of Anisotropy Razors
The qqplot in Figure 5 illustrates the eigenvalue distributions before and after applying these four razors. The first three methods decouple anisotropy while maintaining the linear geometric structure of the data, resulting in eigenvalues that still exhibit distinct partitioning. In contrast, PCS upscales the low-eigenvalue region, ensuring that the corrected compression relies entirely on the contributions of the principal components. The formal method for anisotropy detection involves examining the “self-normalization” property (i.e., $\mathcal{Z}(\mathbf{c})$ tending toward a constant, independent of $\mathbf{c}$mu2018all. Figure 5 illustrates the distribution of $\mathcal{Z}(\mathbf{c})$ before and after applying different anisotropy razors. We observe that remove directions leads to a more dispersed $\mathcal{Z}(\mathbf{c})$ distribution, increasing anisotropy. This occurs because truncating certain directions causes the remaining ones to spread more extremely. In contrast, whitening, LW shrinkage, and PCS concentrate the $\mathcal{Z}(\mathbf{c})$ distribution. Notably, PCS achieves more pronounced anisotropy elimination than the other methods by rigidly correcting the eigenvalue distribution.
@@ A.1 Evaluation of Language Models
The evaluation of language models is currently in a state of rapid iterative development, encompassing a variety of tasks, datasets, and benchmarks celikyilmaz2020evaluation; zheng2023judging; tan2024can. Traditional evaluation metrics such as accuracy, F1-score sasaki2007truth, BLEU sellam2020bleurt, and ROUGE lin2004rouge focus on comparing model predictions with annotated labels in downstream tasks. Other metrics like perplexity and cross-entropy loss do not rely on annotated labels and are computed solely based on input text. However, these methods primarily emphasize external evaluation based on model predictions.
Recently, significant efforts have been devoted to exploring the mechanisms by which language models (LMs) process information internally, driving the development of LM self-evaluation wei2024diff; wang2024embedding; wang2024latent; zang2024explanation; dou2024multi independent of specific tasks and model outputs. The concept of “compression as intelligence” has provided an information-theoretic internal evaluation perspective for language models, highlighting that the acquisition of world knowledge by language models is a denoising process sutskever2023compressors; deletanglanguage; wei2024diff; chen2025information. Differential entropy of representations, as a classical information-theoretic measure, effectively quantifies the internal uncertainty of language models cheninside; zhouyin2023understanding. Semantic volume li2025semantic leverages representation-level differential entropy-aware compression metrics to offer a novel perspective for language model evaluation. However, related work has found that such compression can only model the scale of language models and fails to align with their capabilities zang2023improving; wei2024diff; li2025semantic.
We introduce the concept of compression hacking in language model representations, where the noisy directions of LM representations sacrifice spatial uniformity to feign high compression rates. This implies that we can refine the information compression perspective by considering the geometric distortions in the language model’s representation space.
@@ A.2 Anisotropy of Language Models
Anisotropy The anisotropy of language models reflects the geometric properties of the contextual embedding space. Related studies have observed that during sampling, the spatial embeddings of negative samples exhibit anisotropy, which describes how vectors are distributed within the contextual space mimno2017strange; ethayarajh2019contextual. The researchers found that most vectors occupy a relatively narrow cone within the space, and that vectors within this cone tend to have high cosine similarity gaorepresentation. Demeter pointed out that using softmax introduces structural weaknesses in the representation space, leading to bias, a common issue in language models demeter2020stolen. To better quantify the anisotropy of LMs, related work has identified isolated clusters and low-dimensional manifolds in the contextual embedding space, introducing tools for their qualitative and quantitative analysis ethayarajh2019contextual; cai2021isotropy; rudman2022isoscore. However, these tools are mainly based on similarity calculations of embedded representations. What is needed instead is an anisotropy metric that can establish a connection with entropy-based compression metric.
Anisotropy Razors To mitigate anisotropy in language models, existing research has proposed various solutions. Contrastive learning has emerged as a powerful tool for obtaining effective sentence representations, effectively reducing anisotropy by increasing the spatial distance between positive and negative samples gao2021simcse; zhang2022unsupervised; jiang2022promptbert; zang2023extract; zang2024modeling. In this work, we employ post-processing methods applied directly to the representation space as baseline approaches for the anisotropy razor:
  • Remove Directions mu2018all: First, subtract the common mean vector of all word vectors to eliminate global bias; then remove the top high-variance principal component directions via Principal Component Analysis (PCA). This process enhances semantic feature discriminability by eliminating non-semantic common information from word vectors, making the word space distribution more isotropic.
  • Whitening su2021whitening: Zero-center the representations and transform the covariance matrix into an identity matrix, forcing the embedding distribution toward isotropy.
  • LW Shrinkage ledoit2004well: Linearly shrink the sample covariance matrix toward the diagonal matrices to reduce noise interference in high-dimensional data, yielding more stable covariance matrix estimates. This operation mitigates excessive sensitivity in specific directions, promoting isotropic feature distributions.
@@ Appendix B Statistical Properties of Principal Component Smoothing
Lemma B.1 (Asymptotic Optimality of Ledoit-Wolf Shrinkage ledoit2004well).
@@
For our eigenvalue-shrinkage estimator $\hat{\Sigma}_{\text{PCS}}=(1-\beta_{\text{PCS}})\Sigma_{\mathbf{Z}}+\beta_{\text{PCS}}\lambda_{1}\mathbf{I}$, the preserved leading eigenvalue estimation $\lambda_{1}\xrightarrow{p}\lambda_{1}^{*}$ fundamentally alters the bias-variance tradeoff. The bias now originates from minor eigenvalues:
@@ Appendix C Significance Analysis
Figure 7: The correlation coefficients between compression (PCS) and ground truth under different smoothing coefficients.
@@
Figure 8: The Mann-Whitney U tests of compression-anisotropy regression fitting between different models, where *** and **** denote statistical significance at the 0.1% and 0.01% levels respectively.
@@ Appendix D Implementation Details of the Evaluation Pipeline
For the projection dataset, we primarily collected 1,000 data samples from the pretraining corpus (Wiki wikimedia_downloads) and the instruction-tuning dataset (Dolly-15k conover2023free) to derive projection data. By sampling the in-context representations of these data points, we aim to estimate the full model’s representation space, ensuring the convergence of our metrics. Our pipeline defaults to sampling 800 data samples. Figure 9 illustrates the cumulative expected values of different metrics as the number of samples increases. We observe that all metrics converge relatively early to stable values, demonstrating that our refined metrics enable robust evaluation based on the provided projection dataset.